OpenStudy (anonymous):

Please need help! CHECK MY ANSWER! I'll give a medal!! Jacob has money to spend at the Pacific Science Center store! He has a heavy wallet full of 75 coins. There are only quarters and nickels in his wallet, and the total value of the coins is $7.35. Write an equation to represent the situation. Use only one variable. MY ANSWER: $7.35= n + 7 I'm pretty sure it's wrong but please help me to fix it if it is!!

OpenStudy (anonymous):

@acxbox22 @SolomonZelman @Directrix @Nnesha

jimthompson5910 (jim_thompson5910):

"He has a heavy wallet full of 75 coins. There are only quarters and nickels in his wallet" so we know that q+n = 75 where q is the number of quarters and n is the number of nickels. If we solve for q, we get q = 75 - n

jimthompson5910 (jim_thompson5910):

agreed so far?

OpenStudy (anonymous):

yes @jim_thompson5910

jimthompson5910 (jim_thompson5910):

q quarters gives a total value of 0.25q dollars n nickels gives a total value of 0.05n dollars in total, those coins give a total value of 0.25q + 0.05n dollars

jimthompson5910 (jim_thompson5910):

"the total value of the coins is $7.35." so the second equation is 0.25q + 0.05n = 7.35

jimthompson5910 (jim_thompson5910):

to make one equation, you replace q with 75 - n

jimthompson5910 (jim_thompson5910):

since above q = 75 - n

jimthompson5910 (jim_thompson5910):

so what do you get when you do so?

OpenStudy (anonymous):

but what is the value of n?

jimthompson5910 (jim_thompson5910):

it only wants the equation. I don't see it saying "solve for n"

jimthompson5910 (jim_thompson5910):

what do you get when you replace q with 75 - n in 0.25q + 0.05n = 7.35

OpenStudy (anonymous):

0.25 x 75-n + 0.05n?

jimthompson5910 (jim_thompson5910):

you need parenthesis though

jimthompson5910 (jim_thompson5910):

0.25(75 - n) + 0.05n = 7.35

jimthompson5910 (jim_thompson5910):

because 0.25 is multiplying by all of "75-n" and not just the 75

jimthompson5910 (jim_thompson5910):

so if you want to find the value of n, you solve 0.25(75 - n) + 0.05n = 7.35

OpenStudy (anonymous):

Thank you so much @jim_thompson5910

jimthompson5910 (jim_thompson5910):

no problem